OpenStudy (anonymous):

Use two variables and two equations. Solve by any method you choose. Write your answer in the appropriate form. A lady has 11 animals. Some are cats and some are birds. If there are 34 legs among all her animals, how many animals does she have?

OpenStudy (anonymous):

cats=x birds=y total animal = 11 = x+y total legs = 34= 4x+2y if you solve this you can manage to tell how many cats and birds she have?

OpenStudy (anonymous):

eliminate one of the variable and you are good to solve it

OpenStudy (texaschic101):

or you can also use the substitution method

OpenStudy (anonymous):

im lost

OpenStudy (texaschic101):

cats = x birds = y x + y = 11 --->x = 11 - y 4x + 2y = 34 now sub 11 - y in for x in the second equation : 4x + 2y = 34 4(11 - y) + 2y = 34 can you take it from here ?

OpenStudy (anonymous):

kk thank you so much i can do it

OpenStudy (texaschic101):

your welcome :)